What are the solutions to the equation below:
|2x+3|-5=12
I have chosen x=-5 or x=7

Answer by Alan3354(69443)

What are the solutions to the equation below:
|2x+3|-5=12
I have chosen x=-5 or x=7
--------------
Check your answers.
|2x+3|-5=12
x = -5
|-10 + 3 |- 5 =? 12
|-7| - 5 =? 12
7-5 =? 12 Nope.
===============
|2x+3|-5=12
x = 7
|14 + 3| - 5 =? 12
17 - 5 = 12 *** 7 is correct
==============
You don't choose, you figure them out.